Susan recieved a $80.00 gift card. She used it buying coffee that cost $8.18 pre pound. After buying, she had $30.92 left. How many pounds did she buy?
Answer by nerdybill(5404)

Susan recieved a $80.00 gift card. She used it buying coffee that cost $8.18 pre pound. After buying, she had $30.92 left. How many pounds did she buy?
.
Let x = pounds of coffee bought
.
"amt of gift card" - "cost of coffee" = "amt left on gift card"
80 - 8.18x = 30.92
-8.18x = -49.08
x = -49.08/-8.18
x = 6 pounds